OpenStudy (anonymous):

the mean of 4 numbers is 20 if three of the numbers are 18.19, and 20 what is the 4th number

OpenStudy (anonymous):

x+y+z+w/4=20 x=19 y=18 z=20 w=? 18+19+20+w=80 57+w=80 w=23 18+19+20+23/4=20 18+19+20+23=80 CHECK

OpenStudy (anonymous):

For their mean to be 20, their sum must be 80. As the sum of the first three is 67, the fourth must be 23.

OpenStudy (anonymous):

^57